OpenStudy (anonymous):

Which of the following is NOT a Pythagorean triple? 0.5, 1.2, 1.3 6, 8, 10 3, 4, 5 1, 2, 3

OpenStudy (zzr0ck3r):

A) no because they are not integers B) 6^2+8^2 = 36+64=100=10^2 C) 3^2+4^2=9+16=25=5^2 D) 1^2+2^2 = 1+4=5 does not equal 3

OpenStudy (nirmalnema):

the ans is D

OpenStudy (raden):

0.5^2 + 1.2^2 = 1.3^2 @zzr0ck3r

OpenStudy (zzr0ck3r):

A Pythagorean triple consists of three positive integers a, b, and c, such that a2 + b2 = c2.

OpenStudy (raden):

not always be integer :)

OpenStudy (zzr0ck3r):

not only integers satisfy it, but by definition a Pythagorean triple is 3 integers.

OpenStudy (zzr0ck3r):

from what I was taught..

OpenStudy (zzr0ck3r):

pretty sure across the board they are whole numbers.

OpenStudy (raden):

after googled. yeah they are must be integer positive :) http://en.wikipedia.org/wiki/Pythagorean_triple

OpenStudy (zzr0ck3r):

did the Pythagoreans consider lengths of non whole number size?

OpenStudy (zzr0ck3r):

they might have banned them, along with lighting fire next to a mirror, and picking something up if it fell on the ground, or not eating beans.

OpenStudy (anonymous):

D. Because 1+2=3 .the sum of two sides have to be greater than other side.